OpenStudy (anonymous):

If A = James's present age, write an expression for his age seven years from now. A) 7 B) 7A C) 7 - A D) 7 + A @drewbush63 @kakok

OpenStudy (anonymous):

D.

OpenStudy (green_1):

any other questions

OpenStudy (anonymous):

In the expression 2x + 3, the 2 is a __________. A) coefficient B) power C) product D) quotient @green_1

OpenStudy (green_1):

i think its b.)coefficient

OpenStudy (anonymous):

its A.

OpenStudy (anonymous):

Marcie wants to simplify this expression. Which like terms can she combine? 12 - 4x + 3 A) -4x and 3 B) 12 and 3 C) 12 and -4x D) 12, -4x, and 3 @drewbush63

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

If c = 30 and d = 8, then 2c + 4d = A) 512 B) 278 C) 136 D) 92 @drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

Simplify by combining like terms: 3x + 3x + 3x A) 3x3 B) 9x C) 9x2 D) 27x3 @drewbush63

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

B. 9x

OpenStudy (anonymous):

Which expression is equivalent to x + x + x + 5 + 5, no matter what value is substituted in for x? A) 3x + 5 B) 3x + 7 C) x + 10 D) 3x + 10 @drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

Which expression is equivalent to x + x + x + 5 + 5, no matter what value is substituted in for x? A) 3x + 5 B) 3x + 7 C) x + 10 D) 3x + 10 @drewbush63

OpenStudy (anonymous):

D. 3x+10

OpenStudy (anonymous):

Which expression is equivalent to 3(2x +5), no matter what value is substituted in for x? A) 5x + 8 B) 5x + 5 C) 6x + 5 D) 6x + 15 @drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

D. 6x+15

OpenStudy (anonymous):

Solve. 3x = 15 A) x = 4 B) x = 5 C) x = 6 D) x = 12 @drewbush63

OpenStudy (anonymous):

B. x=15

OpenStudy (anonymous):

i mean B. x=5

OpenStudy (anonymous):

Solve. x 2 = 7 A) x = 5 B) x = 7 C) x = 9 D) x = 14 @drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

is that 2x=7?

OpenStudy (anonymous):

Solve. x + 3 = 7 A) x = 2 B) x = 4 C) x = 6 D) x = 8 @drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

@paki Solve. x + 3 = 7 A) x = 2 B) x = 4 C) x = 6 D) x = 8 @swolesammy @drewbush63

OpenStudy (anonymous):

B. x=4

OpenStudy (anonymous):

Jeff volunteers his time by working at an animal shelter. Each year he works for a total of 240 hours. So far this year, he has worked 97 hours. Which equation will solve for how many more hours (h) Jeff will volunteer for? A) 97 = 240 + h B) 97 = 240h C) 240 = 97 + h D) 240 = 97 - h @swolesammy

OpenStudy (anonymous):

@drewbush63

OpenStudy (green_1):

i think it is a

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Jeff volunteers his time by working at an animal shelter. Each year he works for a total of 240 hours. So far this year, he has worked 97 hours. Which equation will solve for how many more hours (h) Jeff will volunteer for? A) 97 = 240 + h B) 97 = 240h C) 240 = 97 + h D) 240 = 97 - h

OpenStudy (rccb):

i think its d because you would subtract the 97 on both sides and that give h=143

OpenStudy (rccb):

or c i meant ^

OpenStudy (anonymous):

A theater sells adult tickets for $8.00 and children's tickets for $5.50. Sally bought some of each kind (not the same number) and paid a total of $78.50. Which equation represents this situation? A) x + y = 78.5 B) 13.5x = 78.5 C) 13.5xy = 78.5 D) 8x + 5.5y = 78.5 @RCCB @drewbush63

OpenStudy (green_1):

i think it is a

OpenStudy (anonymous):

It is given that 2(3 + x) = 6 + 2x. This is an example of the ___________ property. A) associative B) commutative C) distributive D) identity

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

@paki

OpenStudy (rccb):

D was the last one i want to say^^

OpenStudy (green_1):

it is c

OpenStudy (rccb):

c for this one

OpenStudy (anonymous):

Jaylon make twice as much an hour as Sam. Sam makes x dollars an hour. Write an expression describing how much Jaylon makes an hour. A) 2x B) x2 C) 2 + x D) x - 2 @drewbush63 @paki @RCCB

OpenStudy (rccb):

a

OpenStudy (green_1):

the answer is A

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

A number decreased by 23 is 44. Find the number. Which equation can be used to solve this problem? A) x - 23 = 44 B) 23 - x = 44 C) x - 44 = 23 D) 44 - x = 23 @dan815 @paki @RCCB

OpenStudy (rccb):

its A

OpenStudy (green_1):

a

OpenStudy (anonymous):

Miranda wants to buy as many collectible dolls as possible, for $2.50 each. If she has $45.00 to spend, how many dolls can she buy? Which equation BEST represents this situation? A) 45x = 2.5 B) 2.5x = 45 C) x + 2.5 = 45 D) x + 45 = 2.5 @rccb @paki @swolesammy

OpenStudy (rccb):

b

OpenStudy (anonymous):

Convert 3 18 25 to a decimal using long division. A) 0.72 B) 1.38 C) 3.72 D) 3.85 @RCCB

OpenStudy (anonymous):

@sleepyjess @hanner_B_nanner

OpenStudy (anonymous):

Convert 25 18 to a decimal using long division. A) 0.72 B) 1.38 C) 1.72 D) 2.1 @RCCB

OpenStudy (anonymous):

@drewbush63

OpenStudy (anonymous):

@drewbush63

OpenStudy (rccb):

the answer is B

OpenStudy (rccb):

or A

OpenStudy (rccb):

one second

OpenStudy (rccb):

is that the way its written out on the assignment?

OpenStudy (anonymous):

Tim counted 24 mole rats in an exhibit at the zoo. Then he realized that he forgot to count x mole rats. A. Write an expression to represent the number of mole rats in the exhibit. Why is this the correct expression? B. Suppose a sign next to the exhibit says there are 36 mole rats. What is the value of x? Explain your reasoning. C. Now that you know there are 36 mole rats in the exhibit, write an expression to represent the number of mole rats that Tim counted. Use x in the expression. @drewbush63 @RCCB CAN SOOME GIVE KE ANSWERS THIS IS FOR REPORT CARD AND DONT WANT TO FAIL

OpenStudy (anonymous):

@RCCB

OpenStudy (green_1):

@ParthKohli

OpenStudy (anonymous):

@RCCB can you pleas ehelp i dont want ot fail this

OpenStudy (rccb):

well b B. 36-24=x 36 is the total amount so you would subtract that from the amount he counted (24) and that would give you the amount he didn't count(x)

OpenStudy (rccb):

A. 24+x=total number of mole rats, he counted 24 so you would add the remaining he didn't count (x) to the 24

OpenStudy (rccb):

your teacher would like B better if you put ---> 24+x=36 you would subtract the 24 from the 36 and that would give x=12 <--- with this explanation

OpenStudy (rccb):

C. 12+x=36, you would take the number he didn't count(12) and subtract that from the total(36) to get x=24 and that is that is how many mole rats Tim counted. idk if you need an explanation for c These are the answers i would use and im good with math 99% of the time so these should be right

OpenStudy (rccb):

A. 24+x=total number of mole rats, he counted 24 so you would add the remaining he didn't count (x) to the 24 B. 24+x=36 you would subtract the 24 from the 36 and that would give x=12 C. 12+x=36, you would take the number he didn't count(12) and subtract that from the total(36) to get x=24 and that is that is how many mole rats Tim counted.

OpenStudy (green_1):

any other questions?